• Main INDEX
  • Monthly INDEX
  • NEXT
    Make New Entry, Make Followup Entry

    User name Kai

    Log entry time 03:29:35 on November 01, 2009

    Entry number 297076


    keyword=existing problems

    1. 103.8K clock on the left arm has double pulsing. The origin of the bad is in the fanout module on second floor. The cable there is messy. So I will ask Ramesh for help to solve it. When you do xscaler, the rate is half of the real case. This is a new problem.

    2. It's my first time happened to be right in front of the left daq when the left tagger is missing. Trace it module by module from upstream of the path, and found that there is a 758 module which is used to adjust the tagger width to be 20ns, on which ,either the tagger is interment conntact or the block itself is poor contact. Just replug it and signal get back. However, this tagger is fooling with me for a long time. I cannot guarantee this is the only problem with the consideration that many people are working there these days which might loose it by accident.

    3. Right arm tagger is unplugged between the patch panel on the detector hut and our DAQ by some person without any notice!!! This in return causes the double pulsing on the left arm tagger which explained why I reported that tagger rate on the left arm is 200hz on the xscaler reading. Remember the first problem

    4. Found the path of bcm u3 which connects between counting house and detector hut is bad. u1,d1,d3 all good. Here "good" means the raw scaler counting on both arms are consistent and agree with the reference one I set up in the counting house. At least currently.

    I hope that everyone who touches those REFERENCE signal such as clock, synchronization, tagger, and bcm, should make a log or an email notice. It is important for analyzing the problem to see if it is a real long existing problem, or just some people change the setup for some test but forget to restore it. It is very time consuming and difficult to trace the problem back. Thanks.

    A copy of this log entry has been emailed to: xiaochao,rom,dwang,rsubedi,deng,pkpan